# MSGNav: Unleashing the Power of Multi-modal 3D Scene Graph for Zero-Shot Embodied Navigation

### v1

@@ 3.3.2 Adaptive Vocabulary Update (AVU)
Previous approaches achieve open-vocabulary navigation simply by using detectors such as YOLO-W [yolo] and GroundingDINO [groundingdino], but remain constrained by preset vocabulary (e.g., ScanNet-200 [scannet], HM3D [hm3d]). Such constraints hinder VLMs from handling diverse out-of-vocabulary scene representations in the real world. Additionally, preset vocabulary is difficult to accurately describe the image target.
To address this, MSGNav integrates adaptive vocabulary updates during the exploratory reasoning process. We first initialize the vocabulary $V_{t=0}$ with the ScanNet-200. During exploration at time $t$, the VLM inspects edge images $\mathbf{E}_{t}$ and compares objects in $\mathbf{O}_{t}$ to propose novel vocabulary (based on the extension of Eq. 5):
$\mathcal{R}_{t},\hat{V}_{t}=\mathrm{VLM}(\mathbf{S}^{k},\mathbf{F},g,t),$ (6)
Then the novel vocabulary $\hat{V}$ will be incorporated into $V_{t}$: $V_{t}=V_{t-1}\cup\hat{V}_{t}$. By continuously updating $V_{t}$ in this manner, the scene graph progressively captures richer and more accurate object descriptions, supporting navigation beyond preset vocabulary limitations.
@@ 3.3.3 Closed-Loop Reasoning (CLR)
Memory is critical for navigation agents. While prior work focuses mainly on scene memory, other forms of memory can also influence decision-making. In addition to modeling the scene as perception memory, we introduce decision memory $\mathbf{M}$ for closed-loop reasoning. At time step t1t-1, the exploration response $\mathcal{R}_{t-1}$ is stored in a historical action repository $\mathbf{M}_{t}$: $\mathbf{M}_{t}=\mathbf{M}_{t-1}\cup\mathcal{R}_{t},\mathbf{M}_{0}=\varnothing$ . Then $\mathbf{M}_{t}$ is leveraged at time step $t$ to assist subsequent decisions (based on the extension of Eq. 6):
@@
This closed-loop reasoning significantly improved the accuracy of VLM strategy, increasing it from 43.8% to 64.8%.
@@ 3.3.4 Visibility-based Viewpoint Decision (VVD)
Figure 4: Demonstration of the “last-mile” problem. (a) Prior methods select the nearest traversable position after target localization, often failing due to poor or distant viewpoints. (b) Our VVD samples candidate viewpoints and computes visibility, selecting an insightful viewpoint close to the ground-truth (GT) viewpoint for successful navigation.
As a navigation agent, high-level decisions to select a target coordinate, which the low-level planner converts into executable actions. An optimal final decision should choose a viewpoint with clear visibility. Previous zero-shot methods [vlmnav, dynavlm, 3dmem, sg-nav] often assume that knowing the target location is sufficient, and selecting the nearest traversable position as the navigation viewpoint. As shown in Fig. 4(a), this can yield navigation viewpoints with occlusion or poor visibility, causing failure even when the target $\bar{o}$ is correctly localized.
We term this the “last mile” of navigation—identifying an insightful viewpoint despite locating the target $\bar{o}$. To address this, we realize that a good viewpoint must have clear visibility of the target. To achieve this, we propose a Visibility-based Viewpoint Decision (VVD) module (Algo. 2). As shown in Fig. 4(b), VVD uniformly samples candidate viewpoints around target $\bar{o}$ within radii $\mathbf{R}=\{r_{j}\}_{j=1}^{N_{R}}$. For each candidate $\mathbf{v}_{i}$, VVD computes a visibility score by evaluating occlusion between $\mathbf{v}_{i}$ and points in the target point cloud $\mathcal{PC}_{\bar{o}}$. The viewpoint with the highest visibility score $\textbf{v}_{best}$ is selected as the navigation goal. Compared with traditional nearest-point selection, VVD consistently yields viewpoints with superior target visibility.
@@
1:Target object $\bar{o}$, Scene point cloud $\mathcal{PC}$, Radii set $\mathbf{R}$, Samples number $K$, Camera height $h,$
@@ 4.1 Experimental Setting
We evaluate on two goal-oriented open-vocabulary navigation benchmarks: GOAT-Bench [goatbench] (multi-modal lifelong open-vocabulary navigation), HM3D-OVON [ovon] (open-vocabulary navigation). The metric of navigation task includes Success Rate ($\text{SR}=\frac{N_{\text{success}}}{N_{\text{total}}}$) and Success Path Length ($\mathrm{SPL}=\frac{1}{N_{\text{total}}}\sum_{i=1}^{N_{\text{total}}}S_{i}\frac{l^{s}_{i}}{\max(l^{s}_{i},l^{a}_{i})}$), where $S_{i}\in\{0,1\}$ indicates success of task, $l^{s}_{i}$ and $l^{a}_{i}$ represent shortest path length and agent path length, respectively. To demonstrate generalization capabilities, we evaluate approaches on the “Val Unseen” set of GOAT-Bench and HM3D-OVON datasets. We adopt GPT-4o (2024-08-06 version) as the choice of VLM by directly utilizing the OpenAI API. We also provide results using Qwen-VL-Max (2025-08-13 version) Aliyun API in the Appendix.
@@ 4.2.1 HM3D-OVON
Table 1: Experiments on the ”Val Unseen” split of HM3D-OVON.
Table 1 demonstrates that our proposed MSGNav significantly outperforms all other methods in both SR and SPL settings. Notably, while achieving the highest SR of 48.3%, MSGNav also exhibits an exceptionally high SPL, surpassing the previous best method MTU3D [mtu3d] by 7.5% in SR and 14.9% in SPL. This shows that our approach effectively balances success rate and efficiency in the navigation exploration process. This highlights its robust generalization capability for unseen scenarios.
@@ 4.2.2 Goat-Bench
Table 2: Experiments on the ”Val Unseen” split of GOAT-Bench. ”†” denotes the results we reproduced due to different settings.
@@
M3DSG VVD AVU CLV Overall Category Language Image Overall Category Language Image
@@
Table 3: Component ablation experiment across the first episode of each scene on the ”Val Unseen” split of GOAT-Bench. The first row without any module, which represents our baseline model 3D-Mem [3dmem] results. ”VVD”, ”AVU”, and ”CRV” represent the Visibility-based Viewpoint Decision module, Adaptive Vocabulary Update module, and Closed-loop Reasoning and Verification module.
The results in Table 2 highlight the outstanding performance of our MSGNav when tackling multimodal targets, lifelong navigation, and open-vocabulary task settings. 29.6% SPL metrics compared to both previously training-based and train-free methods. It surpasses the previous state-of-the-art training-based method MTU3D [mtu3d] 4.8% SR and 2.1% SPL, demonstrating its superior navigation capabilities. These results highlight the effectiveness of our multi-modal scene graph M3DSG for multi-modal open-vocabulary navigation tasks.
@@
Table 4: Scene graph experiment across the first episode of each scene on the ”Val Unseen” split of GOAT-Bench. ”Node-only” indicates Concept-graph [Conceptgraphs] without object relation edges.
@@ 4.3 Ablation Analysis
We will analyze the effectiveness and advancement of each module in the MSGNav system. Specifically, this includes the effectiveness of each component on MSGNav; the advantage of M3DSG’s multimodal edges over traditional scene graph edges; and demonstrating how the VVD module aids in “last-mile” decision-making.
@@ 4.3.1 Effective of Component
Table 3 demonstrates the impact of each component. Introducing MSG3D first yielded a 15.0% improvement in SR and a 7.8% increase in SPL (row 2). Subsequently, VVD further delivered a comprehensive boost of 12.5% in SR and 6.7% in SPL (row 3). Notably, introducing either AVU (row 4) or CLV (row 5) alone only improves certain categories. This is because AVU’s new vocabulary may introduce perceptual attribute bias, while CLV may hinder navigation toward language-described targets. Simultaneously incorporating AVU and CLV (row 6) creates a complementary effect, achieving optimal performance. These results validate the effectiveness of each component.
@@ 4.3.2 Advantage of M3DSG
Table 4 shows the experiment results of comparing MSGNav using Node‑only, Concept‑graph [Conceptgraphs], and our M3DSG (). Concept‑graph outperforms Node‑only by 4.4% SR and 1.5% SPL, especially for Language and Image goals, underscoring the value of relationship edges. M3DSG further boosts these categories with a total 3.8% SR and 4.3% SPL, showing its effectiveness in scene understanding. It is worth noting that node-only performs well on category goals, likely because such tasks rely on concise object sets rather than complex contextual information.
@@
Table 5: Experiments on the ”Val Unseen” split of GOAT-Bench with various success thresholds. VVD module represents the Visibility-based Viewpoint Decision module.
4.3.3 Decision-making for ”Last-mile”
Table 5 validates the effectiveness of VVD in addressing the “last‑mile” problem. Under the 0.25 m threshold, removing VVD module yields only a 33.91% success rate. Relaxing the threshold significantly increases success, indicating that failures often occur when stopping close to but misaligned with the optimal viewpoint due to incorrect navigation point selection. With the VVD module, SR improves to 51.91%, primarily recovering cases within 0.25–1 m.
Fig. 5 further confirms the reliability of the candidate viewpoint score estimated by the VVD module: higher scores correlate with smaller distances to the GT viewpoints, and scores above 0.6 consistently correspond to distances below the success threshold. Nevertheless, Table 5 also shows gains under relaxed thresholds, suggesting VVD mitigates but does not entirely eliminate the “last‑mile”.
@@ 5 Conclusion
In this paper, we propose MSGNav, a zero‑shot embodied navigation framework built upon a Multi-modal 3D Scene Graph (M3DSG) that preserves visual information for efficient construction, robust perception, and unconstrained vocabulary. Meanwhile, we integrate a visibility‑based viewpoint decision module in MSGNav to resolve the last‑mile problem. Sufficient experiments on GOAT‑Bench and HM3D‑OVON demonstrate consistent and notable gains over existing methods, highlighting the effectiveness of multi‑modal scene representations in embodied navigation.
Limitations and future work. Despite these advantages of MSGNav, scene graph-based methods still face low inference efficiency due to the latency of VFMs and VLMs, suggesting future work on faster graph construction and inference for real‑time deployment. Additionally, while the “last-mile” problem has been mitigated by the VVD module, it remains unresolved. Exploring reinforcement learning approaches, such as active target recognition, to further solve this challenge is also a concern worthy of attention.

### v2

@@ 3.3.2 Adaptive Vocabulary Update (AVU)
Existing methods achieve open-vocabulary navigation simply by using detectors, e.g., YOLO-W [yolo] and GroundingDINO [groundingdino], but remain constrained by preset vocabulary (e.g., ScanNet-200 [scannet], HM3D [hm3d]). Such constraints hinder VLMs from handling diverse out-of-vocabulary scene representations in the real world. Additionally, preset vocabulary is difficult to accurately describe the image target.
To address this issue, MSGNav integrates adaptive vocabulary updates with preserved visual evidence during the exploratory reasoning process. We first initialize the vocabulary $V_{t=0}$ with the ScanNet-200. During exploration at time $t$, the VLM inspects edge images $\mathbf{E}_{t}$ and compares objects in $\mathbf{O}_{t}$ to propose required novel vocabulary $\hat{V}_{t}$ (based on the extension of Eq. 5):
$\mathcal{R}_{t},\hat{V}_{t}=\mathrm{VLM}(\mathbf{S}^{k},\mathbf{F},g,t).$ (6)
Then the novel vocabulary $\hat{V}$ will be incorporated into $V_{t}$: $V_{t}=V_{t-1}\cup\hat{V}_{t}$. By continuously updating vocabulary $V_{t}$ in this manner, the scene graph progressively captures richer and more accurate object descriptions, supporting navigation beyond preset vocabulary limitations.
@@ 3.3.3 Closed-Loop Reasoning (CLR)
Memory is critical for navigation agents. While prior work focuses mainly on scene memory, other forms of memory can also influence decision-making. In addition to modeling the scene as perception memory, we introduce the decision memory $\mathbf{M}$ for closed-loop reasoning. At time step t1t-1, the exploration response $\mathcal{R}_{t-1}$ is stored in a historical action repository $\mathbf{M}_{t}$ as $\mathbf{M}_{t}=\mathbf{M}_{t-1}\cup\mathcal{R}_{t},\mathbf{M}_{0}=\varnothing$ . Then $\mathbf{M}_{t}$ is utilized at time step $t$ to assist subsequent decisions (based on the extension of Eq. 6) as:
@@
This closed-loop reasoning enhances the accuracy of VLM strategies by memorizing past decision feedback to inform current decisions.
@@ 3.3.4 Visibility-based Viewpoint Decision (VVD)
Figure 4: Demonstration of the “last-mile” problem. (a) Previous methods select the nearest traversable position after target localization, and often fail due to poor or distant viewpoints. (b) Our VVD samples candidate viewpoints and computes visibility, and is able to select a suitable viewpoint close to the ground-truth (GT) viewpoint for successful navigation.
For a navigation agent, its high-level decision process involves the selection of a target coordinate, while the low-level planner converts it into executable actions. Therefore, an optimal final decision should choose a viewpoint with clear visibility to facilitate subsequent operations. Previous zero-shot methods [vlmnav, dynavlm, 3dmem, sg-nav] often assume that knowing the target location is enough, and select the nearest traversable position as the navigation viewpoint. However, as shown in Fig. 4(a), this may yield navigation viewpoints with occlusion or poor visibility, causing failure even when the target $\bar{o}$ is correctly localized.
We term this the “last-mile” problem of navigation—identifying a suitable viewpoint despite correctly locating the target $\bar{o}$. To solve this problem, we realize that a good viewpoint should have clear visibility of the target. To achieve this goal, we propose a Visibility-based Viewpoint Decision (VVD) module (in Algorithm 2). As shown in Fig. 4(b), VVD uniformly samples candidate viewpoints around target $\bar{o}$ within radii $\mathbf{R}=\{r_{j}\}_{j=1}^{N_{R}}$. For each candidate $\mathbf{v}_{i}$, VVD computes a visibility score by evaluating occlusion between $\mathbf{v}_{i}$ and target point cloud $\mathcal{PC}_{\bar{o}}$. The viewpoint with the highest visibility score $\textbf{v}_{best}$ is selected as the navigation coordinate. Compared with traditional nearest-point choice, VVD consistently yields viewpoints with superior target visibility.
@@
1:Target object $\bar{o}$, Scene point cloud $\mathcal{PC}$, Radii set $\mathbf{R}$, Samples number $K$, Camera height $h$, Obstruction distance $\tau$
@@ 4.1 Experimental Setting
We evaluate on two goal-oriented open-vocabulary navigation benchmarks: HM3D-OVON [ovon] (open-vocabulary navigation), and GOAT-Bench [goatbench] (multi-modal lifelong open-vocabulary navigation). The metric of navigation task includes Success Rate ($\text{SR}=\frac{N_{\text{success}}}{N_{\text{total}}}$) and Success Path Length ($\mathrm{SPL}=\frac{1}{N_{\text{total}}}\sum_{i=1}^{N_{\text{total}}}S_{i}\frac{l^{s}_{i}}{\max(l^{s}_{i},l^{a}_{i})}$), where $S_{i}\in\{0,1\}$ indicates success of task, $l^{s}_{i}$ and $l^{a}_{i}$ represent shortest path length and agent path length, respectively. To demonstrate generalization capabilities, we evaluate approaches on the “Val Unseen” set of GOAT-Bench and HM3D-OVON datasets. We adopt GPT-4o (2024-08-06 version) as the choice of VLM by directly utilizing the OpenAI API. We also provide results using Qwen-VL-Max (2025-08-13 version) as the choice of VLM in the Appendix.
@@ 4.2.1 HM3D-OVON
Table 1: Experiments on the “Val Unseen” split of HM3D-OVON.
Table 1 demonstrates that our proposed MSGNav significantly outperforms all other methods on both SR and SPL metrics in open-vocabulary navigation. Notably, while achieving the highest SR of 48.3%, MSGNav also exhibits an exceptionally high SPL, surpassing the previous state-of-the-art training-based method MTU3D [mtu3d] by 7.5% in SR and 14.9% in SPL. This shows that our approach effectively balances accuracy and efficiency in the navigation exploration process for unseen scenarios.
@@ 4.2.2 Goat-Bench
The results in Table 2 also show the outstanding performance of our MSGNav in tackling multimodal goals, lifelong navigation task settings. MSGNav achieves the best performance with 52.0% SPL and 29.6% SPL compared to both previously training-based and train-free methods. It surpasses the previous state-of-the-art training-based method MTU3D [mtu3d] by 4.8% SR and 2.1% SPL, demonstrating its superior navigation capabilities. These results highlight the effectiveness of our multi-modal scene graph in tackling multi-modal lifelong navigation tasks.
@@
Table 2: Experiments on the “Val Unseen” split of GOAT-Bench. “†” denotes the results we reproduced due to different settings.
@@
M3DSG VVD AVU CLR Overall Category Language Image Overall Category Language Image
@@
Table 3: Component ablation experiment across the first episode of each scene on the “Val Unseen” split of GOAT-Bench. The first row without any module, which represents our baseline model 3D-Mem [3dmem] results. “VVD”, “AVU”, and “CRV” represent the Visibility-based Viewpoint Decision module, Adaptive Vocabulary Update module, and Closed-loop Reasoning and Verification module.
@@
Table 4: Scene graph experiment across the first episode of each scene on the “Val Unseen” split of GOAT-Bench. “Node-only” indicates Concept-graph [Conceptgraphs] without object relation edges. “Traditional graph” indicates Concept-graph [Conceptgraphs].
@@ 4.3 Ablation Analysis
In this section, we further analyze the effectiveness and advancement of each module in the MSGNav system. Specifically, this includes component ablation, the advantages of multimodal edges, and demonstrating how the VVD module aids in “last-mile” decision-making.
@@ 4.3.1 Effective of Component
Table 3 demonstrates the impact of each component. Firstly, introducing MSG3D yielded an improvement of 15.0% in SR and 7.8% in SPL (row 2). Subsequently, VVD further gains 12.5% in SR and 6.7% in SPL (row 3). Notably, introducing either AVU (row 4) or CLR (row 5) alone can only improve a part of categories. This is because AVU’s new vocabulary may introduce some insufficient perception results, while the strict decision-making of CLR may hinder navigation toward detailed language-described targets. However, the optimal performance of simultaneously incorporating AVU and CLR (row 6) shows that AVU and CLR are complementary. AVU can supplement extra perception information for the strict decision-making of CLR, while CLR can address the insufficient perception of AVU. These results validate the effectiveness of each component.
@@ 4.3.2 Advantage of M3DSG
Table 4 shows the experimental results comparing MSGNav using Node‑only, Concept‑graph [Conceptgraphs], and our M3DSG. Concept‑graph outperforms Node‑only by 4.4% in SR and 1.5% in SPL, especially for Language and Image goals, underscoring the value of relationship edges. M3DSG further boosts these categories with a total of 3.8% SR and 4.3% SPL, showing its effectiveness in scene understanding. It is worth noting that node-only performs well on category goals, mainly because such tasks rely on concise object sets rather than complex contextual information.
@@
Table 5: Experiments on the “Val Unseen” split of GOAT-Bench with various success thresholds. VVD module represents the Visibility-based Viewpoint Decision module.
4.3.3 Decision-making for “Last-mile”
Table 5 validates the effectiveness of VVD in addressing the “last‑mile” problem. Navigating without VVD yields only a 33.91% SR at the standard threshold of $d$ = 0.25 m. As the threshold is relaxed, its SR can be significantly improved, indicating that lots of failed tasks stop near the target due to incorrect navigation point selection. Navigating with VVD can improve SR to 51.91%, which primarily recovers cases within 0.25–1 m. Fig. 5 further confirms the reliability of the visibility score estimated by the VVD: the higher-scoring viewpoint is closer to the GT viewpoint, and viewpoints with scores above 0.6 are always sufficiently close to the GT viewpoints. Nevertheless, Table 5 also shows that the gains of navigating with VVD can still be achieved as the threshold is relaxed, demonstrating that VVD mitigates but does not eliminate the “last‑mile” problem.
@@ 5 Conclusion
In this paper, we propose the MSGNav, a zero‑shot embodied navigation framework built upon a Multi-modal 3D Scene Graph (M3DSG) that preserves visual information for efficient construction, robust perception, and unconstrained vocabulary. Meanwhile, we integrate a visibility‑based viewpoint decision module in MSGNav to resolve the last‑mile problem. Extensive experiments on GOAT‑Bench and HM3D‑OVON datasets demonstrate consistent and notable gains over existing methods, highlighting the effectiveness of the proposed multi‑modal scene representations in embodied navigation.
Limitations and future work. Despite these advantages of MSGNav, scene graph-based methods still face low inference efficiency due to the latency of VFMs and VLMs, suggesting future work on faster graph construction and inference for real‑time deployment. Additionally, while the “last-mile” problem has been mitigated by the VVD module, it is not fully resolved. Exploring reinforcement learning approaches, e.g. active target recognition, to further solve this challenge is also a way worthy of attention.